Grade 12Algebra

please please answer it fast........
if 3 and 5 be the distance between the parallel sides and 30 degree is the angle between two adjecent side of a parallelogram then its area is???
a question from co-ordinate geometry

Mrunal Sonawane
7 Years agoGrade 12

1 Answer

Arun
ApprovedApproved Tutor Answer7 Years ago
Dear student
 
Area of Parallelogram = p1 * p2 / sinx = 3 * 5 / sin 30 = 20 square unit
 
hope it helps
 
Regards
Arun (askIITians forum expert)